OpenStudy (anonymous):

Identify a line that will be parallel and perpendicular to the given line: y= -2x- 7

OpenStudy (mathstudent55):

A line can't be both parallel and perpendicular to another line.

OpenStudy (anonymous):

no, like one of each

OpenStudy (xapproachesinfinity):

your problem is set incorrectly check the original problem and fix your post

OpenStudy (mathstudent55):

Are there choices given?

OpenStudy (mathstudent55):

Parallel lines have the same slope. In the equation y = mx + b, m is the slope.

OpenStudy (mathstudent55):

In your case, you have line y =-2x - 7 Compare that with y = mx + b What is the slope of your line?

OpenStudy (anonymous):

No they just asked me to identify and parallel and perpendicular line for y= -2x-7

OpenStudy (mathstudent55):